# photon/.style=decorate, decoration=snake, draw=black

### v1

@@ III Quantum Deep Neural Networks
QDNNs integrate principles of quantum mechanics into deep learning architectures, offering a novel computational framework distinct from classical deep neural networks (CDNNs). While CDNNs process data using weighted sums and nonlinear activation functions in Euclidean space, QDNNs manipulate quantum states in Hilbert space via unitary transformations. PennyLane, a quantum computing library, enables the development of these models by integrating quantum circuits—either simulated or executed on quantum hardware—into classical machine learning frameworks. Although the long-term goal is to deploy QDNNs on fault-tolerant quantum processors, current implementations primarily rely on simulators, which serve as an efficient platform for testing and optimizing quantum neural architectures. These approaches offer alternative algorithmic methods that, in some cases, complement classical deep learning.
@@
$|\psi^{\prime}\rangle=U|\psi\rangle.$
@@
$R_{x}(\theta)=e^{-i\theta X/2}=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix}.$
@@
$R_{y}(\theta)=e^{-i\theta Y/2},\quad R_{z}(\theta)=e^{-i\theta Z/2}.$
@@
$|\psi_{\text{final}}\rangle=U_{L}U_{L-1}\cdots U_{1}|0\rangle^{\otimes n}.$
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$
@@
The final step in a QDNN is measurement, which extracts classical information from the quantum state. In quantum mechanics, measurements correspond to taking the expectation values of Hermitian observables. The expectation value of an observable $O$ for a quantum state $|\psi\rangle$ is computed as
$\langle O\rangle=\bra{\psi}O\ket{\psi}.$
@@
$\langle Z\rangle=\langle\psi|Z|\psi\rangle=|\langle 0|\psi\rangle|^{2}-|\langle 1|\psi\rangle|^{2}.$
@@
$f(\theta)=\langle\psi(\theta)|O|\psi(\theta)\rangle.$
Unlike CDNN which compute gradients via backpropagation, QDNNs employ the parameter shift rule to estimate derivatives of expectation values. This quantum gradient calculation is given by
$\frac{d}{d\theta}\langle Z\rangle=\frac{\langle Z(\theta+\pi/2)\rangle-\langle Z(\theta-\pi/2)\rangle}{2}.$
@@
QDNNs offer a compelling advantage over CDNN due to their ability to exploit the exponential dimensionality of quantum state space. A quantum system with $n$ qubits exists in a Hilbert space of dimension $2^{n}$, allowing QDNNs to represent highly complex transformations with far fewer parameters than classical networks, which scale polynomially. Additionally, quantum entanglement enables QDNNs to model intricate correlations that classical networks struggle to capture efficiently. Superposition further enhances expressivity by allowing quantum states to encode multiple computational pathways simultaneously.
One of the most promising theoretical advantages of QDNNs is their trainability in high-dimensional spaces. CDNNs, particularly in complex optimization landscapes, suffer from barren plateaus, where gradients vanish exponentially as the model scales. Quantum circuits, when properly designed, mitigate this issue through entanglement-assisted learning, preserving trainability across larger architectures. Quantum kernel methods and variational quantum circuits have already demonstrated superior learning efficiency in structured data representations.
@@
As quantum computing technology advances, QDNNs may surpass CDNNs in efficiency and accuracy for specialized tasks. Theoretically, deep quantum circuits with polynomial-depth entangling layers could approximate function classes exponentially faster than classical models. While the realization of quantum advantage in deep learning remains an open question, QDNNs represent a promising avenue for integrating quantum mechanics into artificial intelligence, unlocking new possibilities in optimization, generative modeling, and high-dimensional learning.
The advantage of QDNNs lies in their ability to leverage quantum superposition and entanglement to explore quantum features or function spaces that classical networks may struggle to approximate efficiently. The quantum algorithms themselves may also offer more efficient computation and representation of the complexity of the system. One key advantage is the exponential growth of the Hilbert space, where a quantum system with $n$ qubits exists in a space of dimension $2^{n}$, allowing quantum circuits to represent highly complex transformations with far fewer parameters than classical networks, which scale polynomially. Additionally, quantum circuits can efficiently represent certain functions that classical models struggle to approximate without an exponential number of parameters, particularly those involving highly entangled correlations.
Quantum superposition enables a richer feature representation by allowing multiple computational pathways to be explored simultaneously, while entanglement further enhances the expressivity of quantum circuits by capturing intricate correlations that may be difficult for classical models to learn. This can be particularly beneficial for datasets with hidden structures that quantum circuits can exploit more naturally. Moreover, quantum models have shown potential speedups for specific tasks, such as quantum kernel methods and variational quantum circuits, which have demonstrated the ability to learn complex patterns with fewer computational resources compared to classical methods.
Another important aspect is trainability. CDNNs often suffer from vanishing gradients in high-dimensional parameter spaces, a problem known as barren plateaus. Well-designed quantum circuits, however, can mitigate this issue through entanglement-assisted learning dynamics, making them more efficient in certain scenarios. Despite these advantages, the practical realization of quantum models remains an ongoing challenge due to hardware limitations, and the key is identifying problems where quantum circuits provide a significant edge over classical approaches.
One of PennyLane’s key contributions is enabling researchers to compare QDNNs and CDNNs on the same problems. Since PennyLane allows hybrid quantum-classical models, users can train CDNNs and QDNNs on identical datasets, providing insights into their relative performance. Although quantum supremacy in deep learning is still an open question, early results suggest that QDNNs could offer advantages in feature extraction, kernel-based learning, and generative modeling, particularly for problems with complex high-dimensional structure.
@@
By simulating quantum networks on classical hardware today, PennyLane provides a crucial testbed for developing quantum-enhanced deep learning architectures. As quantum technology progresses, these simulations will transition to real quantum processors, unlocking new possibilities for machine learning and data processing. For now, QDNNs remain an experimental but promising approach, representing a step toward harnessing the full potential of quantum computing for artificial intelligence.
IV Quantum Classification
@@
To better understand the differences in a quantum architecture and a classical one, let us consider a basic classifier DNN. In the simplest case, the architecture can be a basic feedforward single hidden layer classifier designed for binary classification tasks, where the input consists of a set of numerical features, lets chose eight for this example. It operates through two fully connected layers: the first projects the input into an eight-dimensional space and applies a rectified linear activation function to introduce nonlinearity, enhancing its ability to capture complex patterns. The second layer reduces the dimensionality to a single output, applying a sigmoid activation function to constrain the result within a probability range suitable for classification. This type of model learns by adjusting its internal parameters through gradient-based optimization techniques, typically using methods like stochastic gradient descent or adaptive moment estimation, minimizing a binary cross-entropy loss function. Its compact design makes it computationally efficient and effective for structured data classification tasks that require mapping numerical inputs to distinct categories based on learned patterns.
@@
The core computational power of a QDNN lies in its quantum perceptron layers. Each layer applies a unitary transformation $U$ to the quantum state, modifying it while preserving coherence. The evolution of the quantum state follows the transformation, $|\psi^{\prime}\rangle=U|\psi\rangle$.
where $U$ is a unitary matrix satisfying the condition $U^{\dagger}U=I$, ensuring that the transformation remains reversible. Unlike classical networks, which rely on weight matrices and biases, the QDNN utilizes these unitary operations to manipulate quantum states without losing quantum information.
@@
In a multi-qubit system, transformations must be properly structured to maintain coherence and ensure correct processing. Each unitary transformation in a QDNN is tensored with the identity matrix when applied to multiple qubits, ensuring that transformations operate within the appropriate Hilbert space. Mathematically, this is expressed as $U_{\text{layer}}=U_{\text{qubit}}\otimes I$.
This tensoring process ensures that unitary transformations are correctly applied in the quantum circuit without disrupting the overall structure of the network. Additionally, swapped operations are implemented to maintain proper qubit ordering within the layered transformations, further reinforcing the structural integrity of the quantum network.
Once the quantum state has evolved through the network’s hidden layers, the final step is to extract a classification decision from the quantum system. This is achieved through a measurement operation, where an observable $O$, such as the Pauli-Z expectation value, is used to collapse the quantum state into a classical probability. The measurement process follows, $\langle O\rangle=\bra{\psi_{\text{final}}}O\ket{\psi_{\text{final}}}$.
The extracted expectation value serves as the output of the QDNN, which is then mapped to a binary outcome for classification. This final measurement is a critical component of quantum machine learning, as it translates quantum state information back into a form that can be interpreted in the classical world.
QDNNs offer a novel approach to deep learning by harnessing quantum mechanical principles such as unitary evolution, entanglement, and measurement-based classification. While structurally similar to CDNNs in terms of layered transformations, the use of quantum perceptrons, tensoring, and quantum measurement introduces new computational advantages that are not present in classical architectures. As quantum computing technology advances, QDNNs may become a powerful tool for solving problems that are intractable for classical machine learning models.
@@
The way these networks process information also diverges in terms of state evolution. CDNNs compute outputs through matrix-vector multiplication, maintaining a straightforward relationship between inputs and learned weight matrices. QDNNs, however, evolve quantum states via unitary operations, which not only preserve quantum coherence but also allow for quantum superposition and entanglement. Despite these differences, both architectures rely on trainable parameters to optimize performance. In CDNNs, optimization targets the weight matrices $W$ and biases $b$, while QDNNs optimize unitary transformations $U$ using parameterized quantum gates.
Another distinction arises in the methods of gradient computation. CDNNs employ backpropagation, leveraging the chain rule of differentiation to update weights efficiently. QDNNs, on the other hand, use the parameter shift rule, which calculates gradients based on evaluating the function at shifted parameter values $d\langle Z\rangle/d\theta$.
Finally, a critical difference between CDNNs and QDNNs lies in their dimensionality. CDNNs operate in an $n$-dimensional Euclidean space, where input data and transformations are restricted to linear algebra over real or complex numbers. In contrast, QDNNs function within a $2^{n}$-dimensional Hilbert space, providing a vastly richer function representation due to the exponential scaling of quantum state space. This higher-dimensional representation enables QDNNs to potentially capture intricate patterns and correlations more efficiently than their classical counterparts. Despite these fundamental differences, both CDNNs and QDNNs share the goal of learning input-output mappings, albeit within distinct mathematical frameworks.
CDNNs and QDNNs share structural similarities, both consisting of layered transformations, trainable parameters, and optimization via gradient descent. However, QDNNs introduce quantum properties such as entanglement, superposition, and exponentially large state spaces, making them potentially more expressive than CDNNs.
While CDNNs rely on nonlinear activation functions, QDNNs achieve equivalent effects through quantum measurement and entanglement-induced correlations. The primary difference is that classical networks use weight matrices, whereas quantum networks use unitary operators.
Training in QDNNs is fundamentally different due to the unitary constraint and the parameter shift rule, which replaces traditional backpropagation. Additionally, QDNNs encode data differently, using quantum states rather than real-valued vectors, leading to computational advantages for problems with high-dimensional correlations.
Ultimately, the choice between CDNNs and QDNNs depends on the problem domain. While CDNNs are well-suited for classical data, QDNNs may offer superior performance in tasks where quantum correlations and high-dimensional state representations are advantageous, such as quantum chemistry, material science, and certain machine learning tasks that exploit quantum features.
IV.1 Quantum Classification Example
@@
Classification data comes in pairs $\{X,\,y\}$, where $X$ is a vector of $n$ input features and $y$ is the target value 0 or 1. The CDNN used has an architecture of $n$ inputs, a layer of 8 hidden perceptrons, and an output perceptron. A binary cross entropy loss function and the Adam optimization algorithm are used. When classifying, 100 epochs are used for training the CDNN.
@@
In QDNNs, it is common to structure the number of nodes in hidden layers as $\log_{2}(N)$, where $N$ is the number of input variables. This choice is rooted in both quantum computational efficiency and the way quantum systems encode information. In quantum computing, a system with $n$ qubits can represent $2^{n}$ states simultaneously due to quantum superposition. To process $N$ classical input variables efficiently, the natural mapping is to use $\log_{2}(N)$ qubits, ensuring that quantum operations scale optimally without redundancy. Additionally, quantum circuits operate using unitary transformations, and a well-structured quantum neural network should minimize circuit depth. Using $\log_{2}(N)$ nodes in hidden layers maintains an optimal depth-to-width ratio, allowing for efficient parameterization with quantum gates. Many quantum algorithms, such as the Quantum Fourier Transform (QFT) and variational quantum circuits, follow a similar log-scaled structure to ensure logarithmic circuit depth.
@@
Another key reason for this design choice is the role of entanglement in representing feature correlations. Quantum neural networks rely on entanglement to capture dependencies between input variables, and a logarithmic number of nodes allows for efficient encoding and manipulation of these relationships. Using more nodes than necessary introduces redundant parameterization, while too few nodes can limit the network’s ability to learn complex correlations. Additionally, quantum computing offers advantages in compression and speedup compared to classical deep learning architectures. Classical neural networks often require deep layers to capture intricate patterns, but quantum networks, leveraging Hilbert space representations, can encode complex functions with significantly fewer parameters. A logarithmic scaling of hidden nodes ensures efficient quantum learning while maintaining computational feasibility.
@@
Furthermore, in quantum machine learning (QML), feature encoding in a Hilbert space often relies on unitary transformations that act on $\log_{2}(N)$ qubits. Many QML architectures, such as quantum variational classifiers, follow this log-scaled structure to ensure an efficient transformation of input features. This approach aligns with the fundamental principles of quantum computation by optimizing the mapping of classical data into a quantum system, reducing circuit depth, leveraging quantum entanglement, and allowing for quantum speedup. The use of $\log_{2}(N)$ nodes in hidden layers is therefore a natural and effective design choice in quantum neural networks.
@@
The classification pair $\{X,\,y\}$ is converted into the form $\{|\phi^{\mathrm{in}}\rangle,\,|\phi^{\mathrm{target}}\rangle\}$ for the QDNN. This is done by converting an $n$-dimensional input vector $X$ into a $\log_{2}n$-qubit state. Specifically, some input vector
@@
The target value $y=0$ is converted to the target state $|\phi^{\mathrm{target}}\rangle=|0\rangle$, and the target value $y=1$ is converted to the target value $|\phi^{\mathrm{target}}\rangle=|1\rangle$. To mirror the CDNN architecture, the QDNN architecture is set to $\log_{2}n$ input qubits, a layer of 3 hidden quantum perceptrons, and an output quantum perceptron. When classifying, 25 epochs are used for training the QDNN.
Considering output values of 1 to be “positive” and output values of 0 to be “negative,” classification efficiency is measured to be the average of the true positive rate, true negative rate, and accuracy.
(a) CDNN Classification
(b) QDNN Classification
We generate two types of classification data: 1 function data and 3 function data. 1 function signal data ($y=1$) is first generated by sampling each input feature $x_{i}$ from a normal distribution with a mean of 1.0 and a standard deviation of 0.5:
@@
Although QDNNs outperform CDNNs in practically all cases, the factors that seem to play the most critical roles in QDNN outperformance are the number of training pairs, the complexity of the data, and the number of input variables. To optimize data generation for the QDNN outperforming the CDNN, the data complexity and the number of input variables should be high, and the number of training pairs should be low.
When decreasing the number of training pairs, both DNNs decrease in accuracy as expected, but the QDNN outperforms the CDNN significantly more. Considering 3 function data with $0.05\sigma$ noise level, 8 input features, and 500 training pairs yields a classical classification efficiency of 0.6436 and a quantum classification efficiency of 0.8745, but decreasing the number of training pairs to 50 yields a classical classification efficiency of 0.4716 and a quantum classification efficiency of 0.8116. By decreasing the number of training pairs by a factor of 10, the quantum classification efficiency to classical classification efficiency ratio increases by 27%.
When increasing the complexity of the data, both DNNs decrease in accuracy as expected, but the QDNN outperforms more. Considering data with $0.05\sigma$ noise level, 8 input features, and 250 training pairs, using the 1 function generation yields a classical classification efficiency of 0.8378 and a quantum classification efficiency of 0.9678, but increasing to 3 function generation yields a classical classification efficiency of 0.6290 and a quantum classification efficiency of 0.8704. By increasing the complexity of the data by a factor of 3, the quantum classification efficiency to classical classification efficiency ratio increases by 20%.
When increasing the number of input features, the CDNN decreases in accuracy, but the QDNN increases in accuracy. Again, the QDNN consistently outperforms. Considering 3 function data with $0.05\sigma$ noise level, and 250 training pairs, using 8 input features yields a classical classification efficiency of 0.6290 and a quantum classification efficiency of 0.8704, but increasing the number of input features to 16 yields a classical classification efficiency of 0.5327 and a quantum classification efficiency of 0.9359. By increasing the number of variables by a factor of 2, the quantum classification efficiency to classical classification efficiency ratio increases by 27%.
Noise plays a role in QDNN outperformance, but not a significant one. When decreasing the amount of noise in the data, both DNNs increase in accuracy as expected, but the QDNN outperforms more. Considering 1 function data with 8 input features and 250 training pairs data, setting the noise level to $0.2\sigma$ yields a classical classification efficiency of 0.7917 and a quantum classification efficiency of 0.8704, but decreasing the noise level to $0.05\sigma$ yields a classical classification efficiency of 0.8378 and a quantum classification efficiency of 0.9678. By decreasing the amount of noise by a factor of 4, the quantum classification efficiency to classical classification efficiency ratio increases by 5%.

### v2

@@ III Quantum Deep Neural Networks
While the theoretical framework briefly described in the preceding section provides the essential tools and conceptual understanding for characterizing nucleon structure through Compton Form Factors, the practical extraction of these quantities from experimental data presents significant computational challenges. The inherent complexities of the DVCS cross section and the intricate dependencies of CFFs on kinematic variables suggest that a variety of algorithms and approaches should be investigated. With this in mind and to pave the way for the possibility of more precise extractions, this section introduces Quantum Deep Neural Networks (QDNNs). We discuss their fundamental architecture, explaining how these networks leverage quantum mechanical principles such as superposition and entanglement to process information, thereby offering a powerful novel paradigm for complex data analysis.
QDNNs integrate principles of quantum mechanics into deep learning architectures, offering a computational framework distinct from CDNNs. While CDNNs process data using weighted sums and nonlinear activation functions in Euclidean space, QDNNs manipulate quantum states in Hilbert space via unitary transformations. PennyLane [19], a quantum computing library, enables the development of these models by integrating quantum circuits—either simulated or executed on quantum hardware—into classical machine learning frameworks. Although the long-term goal is to deploy QDNNs on fault-tolerant quantum processors, current implementations primarily rely on simulators, which serve as an efficient platform for testing and optimizing quantum neural architectures. These approaches offer alternative algorithmic methods that, in some cases, complement classical deep learning.
@@
$|\psi^{\prime}\rangle=U|\psi\rangle.$ (21)
@@
$R_{x}(\theta)=e^{-i\theta X/2}=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix}.$ (22)
@@
$R_{y}(\theta)=e^{-i\theta Y/2},\quad R_{z}(\theta)=e^{-i\theta Z/2}.$ (23)
@@
$|\psi_{\text{final}}\rangle=U_{L}U_{L-1}\cdots U_{1}|0\rangle^{\otimes n}.$ (24)
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (26)
@@
In this paper, we also frequently use Pennylane’s Strongly Entangling Layers to construct quantum circuits, which are composed of alternating layers of trainable single-qubit rotations and fixed two-qubit entangling gates. Each layer applies parameterized rotations on each qubit followed by a sequence of two-qubit entanglers (e.g., CNOT gates) that connect each qubit $i$ to qubit $(i+r)\bmod M$, where $M$ is the number of qubits and $r$ is a tunable parameter known as the entanglement range. This structure allows the circuit to explore a rich and expressive subspace of the full Hilbert space while maintaining a manageable number of trainable parameters.
The final step in a QDNN is measurement, which extracts classical information from the quantum state. The measurements correspond to taking the expectation values of Hermitian observables. The expectation value of an observable $O$ for a quantum state $|\psi\rangle$ is computed as
$\langle O\rangle=\bra{\psi}O\ket{\psi}.$ (27)
@@
$\langle Z\rangle=\langle\psi|Z|\psi\rangle=|\langle 0|\psi\rangle|^{2}-|\langle 1|\psi\rangle|^{2}.$ (28)
@@
$f(\theta)=\langle\psi(\theta)|O|\psi(\theta)\rangle.$ (29)
Unlike CDNNs, which compute gradients via backpropagation, QDNNs employ the parameter shift rule [20] to estimate derivatives of expectation values. This quantum gradient calculation is given by
$\frac{d}{d\theta}\langle Z\rangle=\frac{\langle Z(\theta+\pi/2)\rangle-\langle Z(\theta-\pi/2)\rangle}{2}.$ (30)
@@
QDNNs offer a compelling advantage over CDNN due to their ability to exploit the exponential dimensionality of quantum state space [21]. A quantum system with $n$ qubits exists in a Hilbert space of dimension $2^{n}$, allowing QDNNs to represent highly complex transformations with far fewer parameters than classical networks, which scale polynomially [22]. Additionally, quantum entanglement enables QDNNs to model intricate correlations that classical networks struggle to capture efficiently [23]. Superposition further enhances expressivity by allowing quantum states to encode multiple computational pathways simultaneously [24].
One of the most promising theoretical advantages of QDNNs is their trainability in high-dimensional spaces [25]. Both CDNNs and QDNNs face optimization challenges in complex landscapes. CDNNs suffer from vanishing gradients [26], where gradients diminish exponentially with network depth, while QDNNs encounter barren plateaus [27], where gradients vanish exponentially as the quantum circuit scales. Quantum circuits, when carefully designed to leverage entanglement, can mitigate barren plateaus, preserving trainability in larger architectures. Quantum kernel methods and variational quantum circuits have demonstrated enhanced learning efficiency for structured data representations compared to classical approaches [23, 28].
@@
As quantum computing technology advances, QDNNs may surpass CDNNs in efficiency and accuracy for specialized tasks. Theoretically, deep quantum circuits with polynomial-depth entangling layers could approximate function classes exponentially faster than classical models [29]. While the realization of quantum advantage in deep learning remains an open question, QDNNs represent a promising avenue for integrating quantum mechanics into artificial intelligence, unlocking new possibilities in optimization, generative modeling, and high-dimensional learning.
@@
Although quantum supremacy in deep learning is still an open question, early results suggest that QDNNs could offer advantages in feature extraction, kernel-based learning, and generative modeling, particularly for problems with complex high-dimensional structure [23, 30, 29, 21, 25]. By simulating quantum networks on classical hardware today, PennyLane provides a crucial testbed for developing quantum-enhanced deep learning architectures. As quantum technology progresses, these simulations will transition to real quantum processors, unlocking new possibilities for machine learning and data processing. For now, QDNNs remain an experimental but promising approach, representing a step toward harnessing the full potential of quantum computing for information extraction in hadronic phenomenology.
In the following subsections, we investigate methods to quantify the quantum advantage of QDNNs over CDNNs in classification and regression tasks using precise performance metrics. Identifying optimal tools for evaluating these metrics is critical, particularly for information extraction in hadronic physics applications.
III.1 Quantum Classification
@@
To better understand the differences between a quantum architecture and a classical one, let us consider a basic classifier DNN. In the simplest case, the architecture can be a basic feedforward single hidden layer classifier designed for binary classification tasks, where the input consists of a set of numerical features. It operates through two fully connected layers: the first projects the input into an eight-dimensional space and applies a rectified linear activation function to introduce nonlinearity, enhancing its ability to capture complex patterns. The second layer reduces the dimensionality to a single output, applying a sigmoid activation function to constrain the result within a probability range suitable for classification. This type of model learns by adjusting its internal parameters through gradient-based optimization techniques, typically using methods like stochastic gradient descent or adaptive moment estimation, minimizing a binary cross-entropy loss function. Its compact design makes it computationally efficient and effective for structured data classification tasks that require mapping numerical inputs to distinct categories based on learned patterns.
@@
The core computational power of a QDNN lies in its quantum perceptron layers. Each layer applies a unitary transformation $U$ to the quantum state, modifying it while preserving coherence. The evolution of the quantum state follows the transformation, $|\psi^{\prime}\rangle=U|\psi\rangle$, where $U$ is a unitary matrix satisfying the condition $U^{\dagger}U=I$, ensuring that the transformation remains reversible. Unlike classical networks, which rely on weight matrices and biases, the QDNN utilizes these unitary operations to manipulate quantum states without losing quantum information.
@@
In a multi-qubit system, transformations must be properly structured to maintain coherence and ensure correct processing. Each unitary transformation in a QDNN is tensored with the identity matrix when applied to multiple qubits, ensuring that transformations operate within the appropriate Hilbert space. Mathematically, this is expressed as $U_{\text{layer}}=U_{\text{qubit}}\otimes I$. This tensoring process ensures that unitary transformations are correctly applied in the quantum circuit without disrupting the overall structure of the network. Additionally, swapped operations are implemented to maintain proper qubit ordering within the layered transformations, further reinforcing the structural integrity of the quantum network.
Once the quantum state has evolved through the network’s hidden layers, the final step is to extract a classification decision from the quantum system. This is achieved through a measurement operation, where an observable $O$, such as the Pauli-Z expectation value, is used to collapse the quantum state into a classical probability. The measurement process follows, $\langle O\rangle=\bra{\psi_{\text{final}}}O\ket{\psi_{\text{final}}}$. The extracted expectation value serves as the output of the QDNN, which is then mapped to a binary outcome for classification. This final measurement is a critical component of quantum machine learning (QML), as it translates quantum state information back into a form that can be interpreted in the classical world.
@@
III.1.1 Quantum Classifier Example
To perform a basic comparison between a CDNN and QDNN classifier, we generate data from two classes and then attempt to separate them. Classification data comes in pairs $\{X,\,y\}$, where $X$ is a vector of $n$ input features and $y$ is the target value 0 or 1. The CDNN used has an architecture of $n$ inputs, a layer of 8 hidden perceptrons, and an output perceptron. A binary cross entropy loss function and the Adam optimization algorithm are used. When classifying, 100 epochs are used for training the CDNN.
In QNNs, the number of nodes in hidden layers is often structured as $\log_{2}(N)$, where $N$ is the number of input variables. This design leverages the principles of quantum computational efficiency and the information-encoding properties of quantum systems. In quantum computing, a system of $n$ qubits can represent $2^{n}$ states simultaneously through quantum superposition and entanglement [31]. To efficiently process $N$ classical input variables, encoding them into $\log_{2}(N)$ qubits provides a compact representation, minimizing redundancy in quantum resources [21].
Because the quantum circuits rely on unitary transformations, and well-designed QNNs aim to minimize circuit depth to reduce computational complexity and mitigate errors in noisy quantum hardware [32]. Structuring hidden layers with $\log_{2}(N)$ nodes helps maintain an optimal balance between circuit depth and width, facilitating efficient parameterization of quantum gates. This logarithmic scaling aligns with several quantum algorithms, such as the Quantum Fourier Transform (QFT) and variational quantum circuits, which often employ log-scaled structures to achieve logarithmic or near-logarithmic circuit depth [33, 34].
Table 1: Comparison of CDNN and QDNN classification efficiency for different data sets. Our default data set is 3 function data with 250 training pairs, 8 input features, and 0.05 noise; each row varies one of these factors while holding the rest at the default. The QDNN classification efficiency is higher than the CDNN classification efficiency for all datasets classified, indicating universal QDNN outperformance in this classification task. Fewer training pairs, higher data complexity, more input features, and lower noise lead to greater QDNN outperformance.
Factor varied Change CDNN eff. QDNN eff. QDNN/CDNN ratio change
Number of training pairs 500 pairs $\rightarrow$ 50 pairs 0.6436 $\rightarrow$ 0.4716 0.8745 $\rightarrow$ 0.8116 27%
Data complexity 1 function data $\rightarrow$ 3 function data 0.8378 $\rightarrow$ 0.6290 0.9678 $\rightarrow$ 0.8704 20%
Number of input features 8 features $\to$ 16 features 0.6290 $\rightarrow$ 0.5327 0.8704 $\rightarrow$ 0.9359 27%
Noise $0.2\sigma$rightarrow $0.05\sigma$ 0.7917 $\rightarrow$ 0.8378 0.8704 $\rightarrow$ 0.9678 5%
@@ III.1.1 Quantum Classifier Example
A key reason for the logarithmic scaling of nodes in QNNs is the role of quantum entanglement in modeling feature correlations. Entanglement enables QNNs to efficiently capture dependencies between input variables, allowing a logarithmic number of qubits to encode and manipulate these relationships effectively [35]. Using more nodes than necessary introduces redundant parameterization, which can increase computational overhead and risk overfitting, while too few nodes may restrict the network’s capacity to learn complex correlations [21]. Furthermore, QNNs offer potential advantages in data compression and computational speedup compared to classical deep learning architectures. CDNNs often require deep architectures with many layers to capture intricate patterns, whereas QNNs, leveraging the high-dimensional Hilbert space, can represent complex functions with significantly fewer parameters [36]. A logarithmic scaling of hidden nodes balances the trade-off between expressive power and computational efficiency, ensuring feasible quantum learning implementations.
@@
In QML, feature encoding into a Hilbert space typically involves unitary transformations applied to a quantum state defined by $\lceil\log_{2}N\rceil$ qubits, where $N$ is the dimension of the input feature space. This logarithmic scaling ensures efficient representation of classical data in a quantum system. Quantum variational classifiers, a common QML architecture, utilize this approach to map input features into a quantum state, optimizing circuit depth and leveraging quantum entanglement for potential computational advantages, such as quantum speedup in specific tasks [37]. The use of $\lceil\log_{2}N\rceil$ qubits in encoding and subsequent layers reflects a design choice that aligns with the principles of quantum computation, balancing expressivity and computational efficiency in quantum neural networks [21].
(a) CDNN Classification
(b) QDNN Classification
@@
The classification pair $\{X,\,y\}$ is converted into the form $\{|\phi^{\mathrm{in}}\rangle,\,|\phi^{\mathrm{target}}\rangle\}$ for the QDNN. This is done by converting an $N$-dimensional input vector $X$ into a $\lceil\log_{2}N\rceil$-qubit state. Specifically, some input vector
@@
The target value $y=0$ is converted to the target state $|\phi^{\mathrm{target}}\rangle=|0\rangle$, and the target value $y=1$ is converted to the target value $|\phi^{\mathrm{target}}\rangle=|1\rangle$. To mirror the CDNN architecture, the QDNN architecture is set to $\lceil\log_{2}N\rceil$ input qubits, a layer of 3 hidden quantum perceptrons, and an output quantum perceptron. When classifying, 25 epochs are used for training the QDNN.
To vary data complexity, we generate two types of classification data: 1 function data, in which the two classes of data are produced using only a single function, and 3 function data, in which the two classes of data are produced using three functions. 1 function signal data ($y=1$) is first generated by sampling each input feature $x_{i}$ from a normal distribution with a mean of 1.0 and a standard deviation of 0.5:
@@
In Table 1, we summarize how our CDNN and QDNN models respond to changes in the number of training pairs, the complexity of the data, the number of input variables, and the noise level. Each of these factors is varied individually, with the other three held constant. We use classification efficiency as the metric to measure the quality of the classifications. Considering output values of 1 to be “positive” and output values of 0 to be “negative,” it is defined to be the average of the true positive (TP) rate, true negative (TN) rate, and accuracy. When comparing the QDNN classification efficiencies with those of the CDNN for any of the datasets, it is clear that the QDNN outperforms the CDNN in the classification of this complex data. The table also indicates that while noise plays a role in QDNN outperformance, the most important factors for QDNN outperformance are the number of training pairs, data complexity, and the number of input variables. Specifically, the QDNN outperformance is higher when the number of training pairs is decreased from 500 to 50, the data complexity is increased from 1 function data to 3 function data, and the number of input variables is increased from 8 features to 16 features.
[…]
$\mathbf{C}^{\text{CDNN}}=\begin{bmatrix}70&24\\ 4&52\end{bmatrix},\quad\mathbf{C}^{\text{QDNN}}=\begin{bmatrix}65&6\\ 9&70\end{bmatrix},$ (43)
where $C_{11}=\text{TP}$, $C_{12}=\text{FP}$ (false positive rate), $C_{21}=\text{FN}$ (false negative rate), and $C_{22}=\text{TN}$. This equates to a 0.8144 classification efficiency for the CDNN compared to a 0.8998 classification efficiency for the QDNN.
Our classifier results suggest a notable advantage in using a QDNN for classification, consistent with previous investigations [39, 40, 41, 42, 43]. However, further investigation is warranted to assess the generalizability of this advantage across diverse datasets and to address potential challenges, such as noise sensitivity and computational overhead, which could impact practical deployment.
III.2 Quantum Regression
QDNNs for regression differ from CDNNs in their approach to modeling continuous outputs, leveraging quantum properties to enhance efficiency in specific scenarios. Unlike CDNNs, which rely on layered architectures with polynomial parameter scaling to fit regression functions, QDNNs encode input data into quantum states within a $2^{n}$-dimensional Hilbert space for $n$ qubits, similar to what was previously described for classification. Theoretically, QDNNs can achieve more compact representations of complex, high-dimensional functions with fewer parameters than CDNNs, particularly for regression tasks involving quantum-native functions, such as those in particle and nuclear physics [29]. For regression tasks, QDNNs use variational quantum circuits (VQCs) to optimize continuous loss functions, such as mean squared error, by adjusting quantum gate parameters via the parameter-shift rule [20]. This allows QDNNs to efficiently capture non-linear relationships in data, particularly for tasks where high-dimensional correlations are prevalent.
A key distinction in regression is the QDNN’s ability to exploit entanglement to model intricate dependencies in continuous data, potentially reducing the training complexity compared to CDNNs, which may require exponentially more parameters to achieve similar accuracy for certain function classes [25]. However, QDNN regression, similar to QDNN classification, faces unique challenges related to barren plateaus that require careful circuit design with controlled entanglement to maintain trainability [28]. While CDNNs are robust for general regression, their performance can degrade in high-dimensional settings due to vanishing gradients [26], whereas QDNNs show promise for specialized regression tasks with structured data.